OpenStudy (sesca89):

WILL FAN AND MEDAL!!! I have a scale of 1/4" = 5'-0", how many centimeters will the leftovers be? What I mean is the feet that can't be evenly put in 5s. PLEASE HELP

OpenStudy (aripotta):

1/4" = 5' just means that 1/4 inch equals 5 feet. so i don't really understand your question?

OpenStudy (sesca89):

Yeah I know, but I just don't know how to draw the other conversions

OpenStudy (sesca89):

Like, the ones that aren't divisible by 5. Like how 3/4" would be 15' is divisible by 5, but 6.5 isn't

OpenStudy (sesca89):

I want to know how I would draw that on paper in centimeters.

OpenStudy (aripotta):

make a proportion?\[\frac{ 0.25 ~inches }{ 5~feet }=\frac{ 6.5~inches }{ x~feet }\]cross-multiply and solve for x

OpenStudy (sesca89):

no, it's 6.5 feet

OpenStudy (sesca89):

i need to know how many cm is 6.5 feet

OpenStudy (aripotta):

ooh ok\[\frac{ 0.25~inches }{ 5~feet }=\frac{ x~inches }{ 6.5~feet }\]

OpenStudy (sesca89):

well...actually i need to know how many cm is in 1.5 feet

OpenStudy (aripotta):

why cm?

OpenStudy (sesca89):

because if it's in 1/4inch scale, then I need to simplify it to cm

OpenStudy (aripotta):

1 inch = 2.54 cm

OpenStudy (sesca89):

wait here I'll show you

OpenStudy (sesca89):